I know headquarters wants us to add that new product line, said Dell Havasi, manager of Billings Companys Office Products1. Compute the Office Products Divisions ROI for this year. 2. Compute the Office Products Divisions ROI for the new produc6. Suppose that the companys minimum required rate of return on operating assets is 14% and that performance is evaluated us
Answer #1

ROI = Operating Income/Average Operating Assets ‘

= 2,347,800/4,000,000

= 58.70%

ROI of new product Line = (9,915,000*35% - 2,607,450)/2,755,000

= 31.32%

ROI for next year = (2,347,800+862,800)/(4,000,000+2,755,000)

= 47.53%

Residual Income = Net Operating income – Average operating Assets*Minimum Required rate of return

This year = 2,347,800 – 4,000,000*14%

=$1,787,800

New product line = 862,800 – 2,755,000*14%

= $477,100

Next year = 1,787,800+477,100

= $2,264,900